OpenStudy (anonymous):

How do I find 85 using parenthesis with the equation 2+3*16+1?

OpenStudy (tkhunny):

There are only a few possibilities. Try some. 2+3*16+1 = 2 + 48 + 1 = 51 -- Nope 2+(3*16)+1 = 2 + 48 + 1 = 51 -- Whoops! Same thing. 2+3*(16+1) = 2 + 3*17 = 2 + 51 = 53 -- Nope. Try other things.

OpenStudy (bishoppatton):

(2+3)*16+1=85

OpenStudy (bishoppatton):

(2+3)*16+1=5*17+=85

OpenStudy (bishoppatton):

@Meia107

OpenStudy (anonymous):

Thanks for your help,but that equals 81.

OpenStudy (anonymous):

Maybe you can help me with this: (w^2-w+1)+(w^3-2w^2+99).

OpenStudy (anonymous):

Also must the sum of 3 polynomials again be a polynomial?

OpenStudy (anonymous):

\[\left( 2+3 \right)*\left( 16+1 \right)=5*17=85\]

OpenStudy (mathmate):

The sum, difference and product of two or more polynomials is always a polynomial. The quotient (division) of two polynomials is a rational function (which \(could\) be a polynomial)

OpenStudy (mathmate):

@meia107 For your problem of adding two polynomials, please read the following link to understand how you can do the addition: http://www.purplemath.com/modules/polyadd.htm

OpenStudy (anonymous):

Thanks for your help. Now can you help me with the other problem